If the perimeter of a certain rectangle is 93 centimeters and the length is 25 centimeters, what is the width?
Found 2 solutions by Maths68, oberobic:
Answer by Maths68(1474)

Perimeter of a rectangle = (2 * length) + (2 * Width)
Given
Perimeter = 93
Length = 25
Width = ?
Perimeter of a rectangle = 2 * length + 2 * Width
93 = (2*25) + 2 * (Width)
93 = 50 + 2 * (Width)
93 - 50=2 * (Width)
43=2 * (Width)
Divide by 2 both sides of the above equation
43/2=2*(Width)/2
Width = 43/2
Width = 21.5


Answer by oberobic(2304)

P = perimeter
P = 2(L+W)
P = 2L + 2W
.
P = 93 cm
.
L = 25 cm
.
93 = 2(25) + 2W
43 = 2W
W = 43/2
.
Done.